OpenStudy (anonymous):

Help plz: Write an equation in point-slope form for the line through the given point with the given slope. (4, - 6); m = 3/5

OpenStudy (amistre64):

point slope form is so named becasue it utilizes a point and the slope in it

OpenStudy (amistre64):

y-Py=m(x-Px) ; where m=slope and P(x,y) attributes to Px and Py

OpenStudy (anonymous):

You lost me on the second one :S

OpenStudy (anonymous):

O snap forgot pic give me a sec plz

OpenStudy (amistre64):

all these forms are reconstructions starting at the idea of finding a slope between 2 points

OpenStudy (amistre64):

the rest of it is just algebraing it into submission

OpenStudy (anonymous):

Here it is

OpenStudy (anonymous):

third one is direct application of point - slope form of a line \[y-y_1=m(x-x_1)\] with your given point and slope

OpenStudy (amistre64):

given 2 points; (x,y) and (Px,Py) we find the slope by subtracting one point from the other as such: ( x, y) -(Px,Py) -------- (x-Px,y-Py) slope(m) is defined as y/x so: \[\frac{y-Py}{x-Px}=m\] by clearing the fraction we obtain the point slope form \[(x-Px)*\frac{y-Py}{x-Px}=m(x-Px)\] \[y-Py=m(x-Px)\]

OpenStudy (amistre64):

now given any slope and a point; we can plug the value into the form: m=3/5 P(4,-6) y-Py = m(x-Px) y--6 = (3/5)(x-4) ; and simplify y+6 = (3/5)(x-4)

OpenStudy (amistre64):

from the point slope form we can algebra our way to the slope intercept form: y-Py = m(x-Px) ; expand y-Py = mx -mPx ; and solve for y y = mx -mPx+Py ; (-mPx+Py) is simply a constant, so to simplify y = mx + C

OpenStudy (anonymous):

Thank you So much man, and thank you to the 10 people who jumped right in to help me :D

OpenStudy (amistre64):

youre welcome